# EnsAI: An Emulator for Atmospheric Chemical Ensembles

### v1

@@ 7 Static Error Covariances
To provide a point of reference for the differences between the GEM-MACH and EnsAI ensemble covariances, we also examine the differences between the GEM-MACH ensemble error covariance and the model error covariance derived in a similar manner to those used operationally at ECCC for surface pollutants (Ménard and Deshaies-Jacques, 2018c; Ménard et al., 2020). ECCC’s operational air quality system produces hourly surface analyses of $\rm{O}_{3}$, $\rm{NO}$, $\rm{NO}_{2}$, $\rm{SO}_{2}$, $\rm{PM}_{2.5}$, $\rm{PM}_{10}$ using error statistics that have a diurnal variation, but are otherwise constant within a two-month period. We refer to these flow-independent error statistics as the ‘static’ error covariances, although this is a bit of a misnomer as they do contain diurnal patterns. Including these comparisons will help quantify the effect of using flow-dependent error covariances over flow-independent error covariances that mirror what is currently used in the operational air quality system at ECCC.
While the ECCC operational error statistics for surface pollutants are formed using the time-average departures of a single forecast from its time-mean and adjusting the error variances using innovation statistics, to isolate the effects of including flow-dependent error covariances, the static error covariances used for this study were formed simply by removing the time-dependence of the GEM-MACH ensemble covariance (but retaining the mean diurnal variation as in the operational system).
@@ 8 Direct Comparisons of the Covariances
Uncertainties on quantities derived from the GEM-MACH ensemble due to the limited size of the ensemble will be estimated via bootstrapping, in which the GEM-MACH ensemble members were resampled (with replacement) 50 times. In all line plots shown below, 1–sigma confidence intervals are illustrated by the lightly-shaded region around the curve. We note that for many of the curves describing the ensemble variance, uncertainties are small enough that the confidence regions are not easily distinguished from the main curve.
@@ 8.1 Covariance Parametrization
[…]
@@
[…]
Defined in this manner, the isotropic correlation length is the expectation value of distance taken with respect to $\varrho$. The factor of $\sqrt{2/\pi}$ in this definition is included so that for a 2D isotropic Gaussian correlation model, $L^{\rm{iso}}$ coincides with the Daley length-scale (Daley, 1993) (as well as the standard deviation of the Gaussian model). We can extend this definition to anisotropic modes $m$ of the correlation length by introducing a multipole expansion of the correlation length as
[…]
where $\phi^{\prime}$ is the angle between $\textbf{{x}}^{\prime}$ and the $x$-axis. With this definition, we have $L[0]=L^{\rm{iso}}$, $L[1]$ is proportional to the dipole moments of the function $\varrho$, and $L[2]$ is related to its quadrupole moments. More details of this expansion can be found in Appendix C.
@@
Figure 4: Examples of correlations and their corresponding $L[1]$ and $L[2]$ values, which are displayed in the upper left corner of each panel. The correlation length modes are taken with respect to the origin. The values for $L[0]$ is the same for all panels and units are selected so that $L[0]=1$.
The set of covariance parameters to compare is then $\{\sigma_{c},L_{ec}[m],L_{cc}[m]\}$, where $L_{ec}[m]$ and $L_{cc}[m]$ are the correlation length modes for the emissions/surface concentration cross-correlation and surface concentration univariate correlations, respectively. Lastly, since the comparisons of $L_{ec}[m]$ and $L_{cc}[m]$ yielded fairly similar conclusions, our analysis will focus on $L_{ec}[m]$ for brevity and only some comparisons for $L_{cc}[m]$ will be displayed.
@@ 8.2 Comparison Statistics
For each covariance parameter $\theta\in\{\sigma_{c},L_{ec},L_{cc}\}$, we will examine the root mean square error $RMSE(\theta)$ and either the correlation coefficient $r(\theta)$ or the anomaly correlation coefficient $ACC(\theta)$, defined by
$RMSE(\theta)=\sqrt{\frac{1}{N}\sum_{i}^{N}|\theta_{i}^{\rm{EnsAI}}-\theta_{i}^{\rm{GM}}|^{2}},$ (9a)
$r(\theta)=\frac{\sum_{i}^{N}\left(\theta_{i}^{\rm{GM}}-\bar{\theta}^{\rm{GM}}\right)\left(\theta_{i}^{\rm{EnsAI}}-\bar{\theta}^{\rm{EnsAI}}\right)}{\sqrt{\sum_{i}^{N}|\theta_{i}^{\rm{GM}}-\bar{\theta}^{\rm{GM}}|^{2}}\sqrt{\sum_{i}^{N}|\theta_{i}^{\rm{EnsAI}}-\bar{\theta}^{\rm{EnsAI}}|^{2}}},$ (9b)
$ACC(\theta)=\frac{\sum_{i}^{N}\left(\theta_{i}^{\rm{GM}}-\theta_{i}^{\rm{c}}-\bar{\theta}^{\rm{GM}}+\bar{\theta}^{\rm{c}}\right)\left(\theta_{i}^{\rm{EnsAI}}-\theta_{i}^{\rm{c}}-\bar{\theta}^{\rm{EnsAI}}+\bar{\theta}^{\rm{c}}\right)}{\sqrt{\sum_{i}^{N}|\theta_{i}^{\rm{GM}}-\theta_{i}^{\rm{c}}-\bar{\theta}^{\rm{GM}}+\bar{\theta}^{\rm{c}}|^{2}}\sqrt{\sum_{i}^{N}|\theta_{i}^{\rm{EnsAI}}-\theta_{i}^{\rm{c}}-\bar{\theta}^{\rm{EnsAI}}+\bar{\theta}^{\rm{c}}|^{2}}}.$ (9c)
In the equations above, the superscripts ‘EnsAI’ and ‘GM’ denote ensemble quantities derived from the EnsAI and GEM-MACH ensembles, respectively, $i$ indexes each quantity (in space and/or time) with a total of $N$ elements, and overbars denote the mean of that quantity. The same statistics will also be computed for the static error covariances to use as a point of reference, which are computed by replacing all variables that have the ‘EnsAI’ superscript in Eqs. (9c) by the analogous quantities for the static error covariances, which will be labeled with the ‘static’ superscript. In Eq. (9c), the superscript ‘c’ refers to a reference value (traditionally a climatological value), which for this work we take as the average value over the whole time period considered (May to August 2016).
@@ 8.3 Comparison Results for Error Variances
We begin with a comparison of the standard deviation of the ammonia surface concentrations $\sigma_{c}$. The error standard deviations as a function of time at the five locations shown in Fig. 1 can be seen in the top panels of Figures 5 to 9. These locations correspond to the places with the highest values in the annual emissions inventory and are ordered with decreasing emissions. From these time series, we can see that EnsAI captures a considerable amount of the temporal variation of the GEM-MACH ensemble variances. There are some notable differences, such as EnsAI’s over-prediction of variance at the location in Iowa in June (Fig. 6). Another significant difference can be seen in Pennsylvania around June 2-3 (Fig. 8), where EnsAI’s variance suddenly spikes. Summary statistics for these time series are shown in tables to the right of each plot. From these tables, we can see that the correlation coefficient $r(\sigma_{c})$ is significantly higher and $RMSE(\sigma_{c})$ is lower for the EnsAI ensemble as compared to the static error covariance for all locations with the exception of Georgia. For the location in Georgia, both $r(\sigma_{c})$ and $RMSE(\sigma_{c})$ are slightly lower for the static error covariance, although the difference in $RMSE(\sigma_{c})$ has low statistical significance.
Figure 5: Time series of components of the error covariances for the location in North Carolina (exact location is shown in Fig. 1) for the GEM-MACH ensemble (blue), the EnsAI ensemble (orange), and the static covariance (green). The top plot shows the time series of the standard deviation of surface ammonia $\sigma_{c}$. Lower plots show time series for the horizontal correlation length modes $L_{ec}[m]$ for wavenumbers $m\leq 2$ for the emissions/surface concentration cross-correlation. Time-mean statistics over the combined time period of July 1-7 and August 1-7 are shown in tables to the right of each plot, where $RMSE$ values in the first and second tables from the top have units of ppbv and km, respectively, and all other table values are unitless.
@@
The largest $RMSE(\sigma_{c})$ values in the model domain generally occur at or near the locations labeled in Fig. 1 (see Figure S2 of the Supplement for $RMSE(\sigma_{c})$ values for whole model domain). The most notable difference between $RMSE(\sigma_{c}^{\rm{EnsAI}})$ and $RMSE(\sigma_{c}^{\rm{static}})$ occurs in North Carolina, where at some locations $RMSE(\sigma_{c}^{\rm{EnsAI}})$ is 3.4 ppbv less than $RMSE(\sigma_{c}^{\rm{static}})$.
The values for $RMSE(\sigma_{c})$ and $ACC(\sigma_{c})$ taken over the whole model domain and evaluation time period are shown in the top row of Table 1, from which we can see that $ACC(\sigma_{c})$ is significantly larger for the EnsAI ensemble as compared to the static covariance, indicating that the EnsAI ensemble was much better at capturing the time variation of $\sigma_{c}^{\rm{GM}}$. The EnsAI ensemble also has a lower $RMSE(\sigma_{c})$ value than the static ensemble.
$ACC$ $RMSE$
EnsAI static EnsAI static
$\sigma_{c}$ $0.73\pm 0.01$ $0.36\pm 0.01$ $0.36\pm 0.01$ ppbv $0.44\pm 0.01$ ppbv
$L_{ec}[0]$ $0.58\pm 0.02$ $0.06\pm 0.02$ $6.51\pm 0.04$ km $9.29\pm 0.03$ km
$L_{ec}[1]/L_{ec}[0]$ $0.87\pm 0.02$ $0.05\pm 0.01$ $0.167\pm 0.001$ $0.346\pm 0.001$
$L_{ec}[2]/L_{cc}[0]$ $0.81\pm 0.03$ $0.05\pm 0.01$ $0.145\pm 0.001$ $0.253\pm 0.001$
$L_{cc}[0]$ $0.58\pm 0.02$ $0.17\pm 0.02$ $5.7\pm 0.3$ km $8.3\pm 0.1$ km
$L_{cc}[1]/L_{cc}[0]$ $0.58\pm 0.02$ $0.06\pm 0.01$ $0.182\pm 0.004$ $0.229\pm 0.002$
$L_{cc}[2]/L_{cc}[0]$ $0.80\pm 0.01$ $0.02\pm 0.01$ $0.174\pm 0.003$ $0.293\pm 0.002$
Table 1: Domain-wide statistics for components of the error covariances for the combined time period of July 1-7 and August 1-7.
@@ 8.4 Comparison Results for Spatial Correlations
As an illustration of the time-dependency of the spatial correlations, Figure 10 shows the emissions/surface concentration spatial correlation $\rho_{ec}$ at the location in North Carolina at three different times, separated by two days and all at 18:00 UTC. Looking at the correlations of the GEM-MACH ensemble in the left column of Fig. 10, we can see that the correlations are often significantly anisotropic and spread out in different directions for different days, reflecting the different meteorological conditions at each time. Comparing this to the correlation in the EnsAI ensemble in the middle column of this figure, it is evident that EnsAI can reasonably reproduce this time-dependence, as the shape of the spatial correlations between the two ensembles look similar. In contrast, the static correlations displayed in the right column of the figure are much closer to being isotropic.
@@
Figure 10: Emissions/surface concentration spatial correlations $\rho_{ec}$ for the GEM-MACH ensemble (left column), the EnsAI ensemble (middle column), and the static covariance (right column) at the location in North Carolina at three different times.
@@
As described in Section 8.1, to make the examination of the correlation matrix more tractable, we will examine the different modes $m$ of the correlation length $L_{ec}[m]$ instead of dealing with the full correlation matrix directly. To illustrate the expansion into modes, Figure 11 shows the correlation length modes $L_{ec}[m]$ as a function of wavenumbers $m$ for the correlations in the top row of Fig. 10. In this example, the isotropic correlation lengths $L_{ec}[0]$ for the different covariances are between 65 and 90 km, while the anisotropic correlation length modes decrease rapidly with increasing wavenumber. Looking at the absolute value of the correlation length modes (top panel of Fig. 11), we can see that the non-isotropic modes for the static covariance is much smaller than that for the GEM-MACH and EnsAI ensembles, indicating less anisotropy in the static correlation. It is also evident from the bottom two panels in this figure that for most modes $L_{ec}[m]$ from the EnsAI ensemble is closer to the GEM-MACH values than the static covariance is. As $L_{ec}[m]$ decreases rapidly with increasing wavenumber, for the remainder of this paper we concentrate mainly on results for modes with $m\leq 2$ for brevity.
@@
Figure 11: Horizontal correlation length modes $L_{ec}[m]$ as a function of wavenumber $m$ for the emissions/surface concentration cross-correlation at the location in North Carolina at 2016-06-01 18:00. The top, middle, and bottom panels show the absolute, real, and imaginary parts, respectively.
@@
Time series of $L_{ec}[m]$ for $m\leq 2$ for the five locations identified in Fig. 1 are shown in the bottom four panels of Figs. 5 to 9. In all of these plots, $L_{ec}[m]$ for the EnsAI ensemble are much closer to the GEM-MACH ensemble value than the static covariance is and captures most of its time variation. While a diurnal pattern is evident in the GEM-MACH ensemble, there is significant non-diural time variation throughout the week that is dependent on meteorological conditions. In Fig. 5, we can see that the EnsAI correlation length modes generally follow the pattern in the original GEM-MACH ensemble, but the results are even more striking for the other four locations (Figs. 6 to 9), where the differences between the EnsAI and GEM-MACH ensembles are often smaller (or not much larger) than the sample variation in the GEM-MACH ensemble, or in other words, the differences between these values are statistically insignificant throughout much of the evaluation time period. Looking at the evaluation statistics taken over the full test time period, which are displayed in the tables on the right side of Figs. 5 to 9, the $RMSE$ values are smaller and correlation coefficients larger for the EnsAI ensemble than for the static covariance for all locations. Many correlation coefficients values for the EnsAI ensemble are above 0.8, while those for the static covariance are mostly below 0.3.
The $RMSE$ values of $L_{ec}[m]$ for the EnsAI ensemble are generally much smaller than those for the static covariance at most grid points (see Figure S3 of the Supplement). In many parts of the model domain the EnsAI ensemble has values $RMSE(L_{ec}[0])\approx 10$ km, $RMSE(L_{ec}[1]/L_{ec}[0])\approx 0.2$, and $RMSE(L_{ec}[2]/L_{ec}[0])\approx 0.1$, while at the same grid points the static covariance has values of $RMSE(L_{ec}[0])\approx 20$ km, $RMSE(L_{ec}[1]/L_{ec}[0])\approx 0.6$, and $RMSE(L_{ec}[2]/L_{ec}[0])\approx 0.5$. For reference, the mean value of $L_{ec}[0]$ over the model domain is roughly equal to 80 km for the GEM-MACH ensemble.
The $ACC$ and $RMSE$ values for both $L_{ec}[m]$ and $L_{cc}[m]$, taken over the whole model domain and evaluation time period, are displayed in the bottom six rows of Table 1. The $RMSE$ values in Table 1 for the EnsAI ensemble are about 3 km, 0.2, and 0.1 lower then those for the static covariance for $L_{ec}[0]$, $L_{ec}[1]/L_{ec}[0]$, and $L_{ec}[2]/L_{ec}[0]$, respectively. The differences in $RMSE$ values for $L_{cc}[m]$ are similar, with the exception of $L_{cc}[1]/L_{cc}[0]$ which is only about 0.05 lower for the EnsAI ensemble. The $ACC$ values for the EnsAI ensemble is significantly higher than for the static covariance for all correlation length modes given in Table 1. The $ACC$ values for the EnsAI ensemble are all between 0.58 and 0.87, while the $ACC$ values for the static covariance are all lower than 0.2. From this we can conclude that the EnsAI ensemble can accurately reproduce the time variations in the horizontal correlations that are not captured by the diurnal pattern of the static covariances.
@@
To get an intuitive understanding of what these differences in correlation length modes correspond to in terms of spatial patterns, Figure 12 shows another set of example correlations that illustrate the deviations in correlation length mode values mentioned in the previous paragraphs. The leftmost panel of Fig. 12 shows an isotropic correlation with $L[0]=80$ km. For illustrative purposes, we will think of this isotropic correlation as a target correlation (i.e. from the GEM-MACH ensemble) and compare it to correlations with different $L[m]$ values to see how different the spatial patterns are. The correlation shown in the panel second from the left has $L[0]$, $L[1]$, and $L[2]$ values that deviate from the isotropic correlation in the leftmost panel by amounts equal to the $RMSE$ values found in Table 1 for EnsAI. This correlation is only slightly anisotropic and slightly smaller than the correlation in the leftmost panel. The correlation in the panel second from the right has values that deviate from the isotropic correlation equal to the $RMSE$ values for the static covariance in Table 1 and is noticeably anisotropic. The correlation in the rightmost panel has values that deviate from the isotropic correlation by amounts equal to the difference between EnsAI and static covariance values that are commonly found in the model domain (Figure S3 of the Supplement), which produces a correlation with a drastically different spatial pattern than the correlation in the leftmost panel.

### v2

@@ 7 Static Error Covariances
To provide a point of reference for the differences between the GEM-MACH and EnsAI ensemble covariances, we also examine the differences between the GEM-MACH ensemble error covariances and the model error covariances derived in a similar manner to those used operationally at ECCC for surface pollutants (Ménard and Deshaies-Jacques, 2018c; Ménard et al., 2020). ECCC’s operational air quality system produces hourly surface analyses of $\rm{O}_{3}$, $\rm{NO}$, $\rm{NO}_{2}$, $\rm{SO}_{2}$, $\rm{PM}_{2.5}$, $\rm{PM}_{10}$ using error statistics that have a diurnal variation, but are otherwise constant within a two-month period. We refer to these flow-independent error statistics as the ‘static’ error covariances, although this is a bit of a misnomer as they do contain diurnal patterns. Including these comparisons will help quantify the effect of using flow-dependent error covariances over flow-independent error covariances that mirror what is currently used in the operational air quality system at ECCC.
While the ECCC operational error statistics for surface pollutants are formed using the time-average departures of a single forecast from its time-mean and adjusting the error variances using innovation statistics, to isolate the effects of including flow-dependent error covariances, the static error covariances used for this study were formed simply by computing the two-month mean of the GEM-MACH ensemble covariance for the training set (done separately for every hour of the day to retaining the mean diurnal variation).
@@ 8 Direct Comparisons of the Covariances
In this section, uncertainties on quantities derived from the GEM-MACH ensemble due to the limited size of the ensemble will be estimated via bootstrapping, in which the GEM-MACH ensemble members were resampled (with replacement) 50 times. In all line plots in this section, 1–sigma confidence intervals are illustrated by the lightly-shaded region around the curve. We note that for many of the curves describing the ensemble variance, uncertainties are small enough that the confidence regions are not easily distinguished from the main curve.
@@ 8.1 Covariance Parametrization
$L^{\rm{iso}}(\textbf{{x}})=\sqrt{\frac{2}{\pi}}\int\varrho(\textbf{{x}},\textbf{{x}}^{\prime})|\textbf{{x}}^{\prime}|d\textbf{{x}}^{\prime},$ (6)
@@
$\varrho(\textbf{{x}},\textbf{{x}}^{\prime})\equiv\frac{\rho(\textbf{{x}},\textbf{{x}}^{\prime})}{\int\rho(\textbf{{x}},\textbf{{x}}^{\prime})d\textbf{{x}}^{\prime}}.$ (7)
Defined in this manner, the isotropic correlation length is the expectation value of distance weighted by $\varrho$. The factor of $\sqrt{2/\pi}$ in this definition is included so that for a 2D isotropic Gaussian correlation model, $L^{\rm{iso}}$ coincides with the Daley length-scale (Daley, 1993) (as well as the standard deviation of the Gaussian model). We can extend this definition to anisotropic modes $m$ of the correlation length by introducing a multipole expansion of the correlation length as
$L[m](\textbf{{x}})\equiv\sqrt{\frac{2}{\pi}}\int e^{im\phi^{\prime}}\varrho(\textbf{{x}},\textbf{{x}}^{\prime})|\textbf{{x}}^{\prime}|d\textbf{{x}}^{\prime},$ (8)
where $\phi^{\prime}$ is the angle between $\textbf{{x}}^{\prime}$ and the $x$-axis. With this definition, we have $L[0]=L^{\rm{iso}}$, $L[1]$ is proportional to the dipole moments of the function $\varrho$, and $L[2]$ is related to its quadrupole moments. More details of this expansion can be found in C.
@@
Figure 4: Examples of correlations and their corresponding $L[1]$ and $L[2]$ values, which are displayed in the upper left corner of each panel. The correlation length modes are taken with respect to the origin. The values for $L[0]$ are the same for all panels and units are selected so that $L[0]=1$.
The set of covariance parameters to compare is then $\{\sigma_{c},L_{ec}[m],L_{cc}[m]\}$, where $L_{ec}[m]$ and $L_{cc}[m]$ are the correlation length modes for the emissions/surface concentration cross-correlation and surface concentration univariate correlations, respectively. Lastly, since the comparisons of $L_{ec}[m]$ and $L_{cc}[m]$ yielded fairly similar conclusions, our analysis will focus on $L_{ec}[m]$ for brevity and only some comparisons for $L_{cc}[m]$ will be discussed.
@@ 8.2 Comparison Statistics
For each covariance parameter $\theta\in\{\sigma_{c},L_{ec},L_{cc}\}$, we will examine the anomaly correlation coefficient $ACC(\theta)$ and the root mean square error $RMSE(\theta)$, defined by
$ACC(\theta)=\frac{\sum_{i}^{N}\left(\theta_{i}^{\rm{GM}}-\theta_{i}^{\rm{c}}-\bar{\theta}^{\rm{GM}}+\bar{\theta}^{\rm{c}}\right)\left(\theta_{i}^{\rm{EnsAI}}-\theta_{i}^{\rm{c}}-\bar{\theta}^{\rm{EnsAI}}+\bar{\theta}^{\rm{c}}\right)}{\sqrt{\sum_{i}^{N}|\theta_{i}^{\rm{GM}}-\theta_{i}^{\rm{c}}-\bar{\theta}^{\rm{GM}}+\bar{\theta}^{\rm{c}}|^{2}}\sqrt{\sum_{i}^{N}|\theta_{i}^{\rm{EnsAI}}-\theta_{i}^{\rm{c}}-\bar{\theta}^{\rm{EnsAI}}+\bar{\theta}^{\rm{c}}|^{2}}},$ (9a)
$RMSE(\theta)=\sqrt{\frac{1}{N}\sum_{i}^{N}|\theta_{i}^{\rm{EnsAI}}-\theta_{i}^{\rm{GM}}|^{2}}.$ (9b)
In the equations above, the superscripts ‘EnsAI’ and ‘GM’ denote ensemble quantities derived from the EnsAI and GEM-MACH ensembles, respectively, $i$ indexes each quantity (in space and/or time) with a total of $N$ elements, and overbars denote the mean of that quantity. The same statistics will also be computed for the static error covariances to use as a point of reference, which are computed by replacing all variables that have the ‘EnsAI’ superscript in Eqs. (9b) by the analogous quantities for the static error covariances, which will be labeled with the ‘static’ superscript. In Eq. (9a), the superscript ‘c’ refers to a reference value (traditionally a climatological value), which for this work we take as the two-month mean value derived from the GEM-MACH ensemble for 2016.
@@ 8.3 Comparison Results for Error Variances
We begin with a comparison of the standard deviation of the ammonia surface concentrations $\sigma_{c}$. The top panels of Figures 5 to 9 show time series of $\sigma_{c}$ for the month of June 2015 at the five locations shown in Fig. 1. These locations correspond to the places with the highest values in the ammonia emissions inventory and are ordered with decreasing emissions. As seen in these panels, $\sigma_{c}$ for the original GEM-MACH ensemble (shown with the blue curve) displays sporadic spikes associated with changing weather conditions, many of which can also be seen in $\sigma_{c}$ from the EnsAI ensemble (shown with the orange curve) but not for the static error covariance (shown with the green curve). In these time series, we can see that EnsAI captures a considerable amount of the temporal variation of the GEM-MACH ensemble variances. There are some notable differences, such as EnsAI’s over-prediction of variance at the location in Ohio on June 4-5 (Fig. 6) and in Pennsylvania on June 6 (Fig. 8) and under-prediction in Georgia on June 23, but for the most part EnsAI produces values for $\sigma_{c}$ that are closer to the GEM-MACH ensemble value than that derived from the static error covariances.
Figure 5: Time series of components of the error covariances for the location in North Carolina (exact location shown in Fig. 1) for the GEM-MACH ensemble (blue), the EnsAI ensemble (orange), and the static covariance (green) for June 2015. The top plot shows the time series of the standard deviation of surface ammonia $\sigma_{c}$. Lower plots show time series for the horizontal correlation length modes $L_{ec}[m\leq 2]$ for the emissions/surface concentration cross-correlation.
@@
The top panel of Figure 10, which shows $ACC(\sigma_{c})$ throughout the test period, show dramatic differences between EnsAI and the static covariances, with EnsAI having larger $ACC$ values for all months. EnsAI has $ACC(\sigma_{c})$ values around 0.8 for most months (excluding December and January), while the static covariances have values between 0.2 and 0.4 for February to November and values not far from zero for December and January, indicating that the EnsAI ensemble was much better at capturing the time variation of $\sigma_{c}$ in the GEM-MACH ensemble.
The $RMSE$ values for $\sigma_{c}$ over the model domain for each month in 2015 are shown in the top panel of Figure 11. The $RMSE(\sigma_{c})$ values for EnsAI are lower than the static covariance values for all months, with both EnsAI and the static covariances performing poorer in December and January as compared to the rest of the year. While the differences in $RMSE$ values between EnsAI and the static covariances taken over the whole model domain are less than a ppbv (see Figure S2 of the Supplement for a comparison to the mean values of $\sigma_{c}$), the $RMSE(\sigma_{c})$ values near locations with high emissions can vary significantly. For instance, in the summer, the $RMSE(\sigma_{c})$ for the static covariances are 4 ppbv higher at the location in North Carolina than for EnsAI (see Figure S3 of the Supplement).
Figure 10: ACCACC values for $\sigma_{c}$, $L_{ec}[0]$, $L_{ec}[1]/L_{ec}[0]$, and $L_{ec}[2]/L_{ec}[0]$ over the model domain for each month in 2015.
Figure 11: RMSERMSE values for $\sigma_{c}$, $L_{ec}[0]$, $L_{ec}[1]/L_{ec}[0]$, and $L_{ec}[2]/L_{ec}[0]$ over the model domain for each month in 2015.
@@ 8.4 Comparison Results for Spatial Correlations
As an illustration of the time-dependency of the spatial correlations, Figure 12 shows the emissions/surface concentration spatial correlation $\rho_{ec}$ at the location in North Carolina at three different times, separated by two days and all at 18:00 UTC. Looking at the correlations of the GEM-MACH ensemble in the left column of Fig. 12, we can see that the correlations are often significantly anisotropic and spread out in different directions for different days, reflecting the different meteorological conditions at each time. Comparing this to the correlation in the EnsAI ensemble in the middle column of this figure, it is evident that EnsAI can reasonably reproduce this time-dependence, as the shape of the spatial correlations between the two ensembles look similar. In contrast, the static correlations displayed in the right column of the figure are much closer to being isotropic (at all times).
@@
Figure 12: Emissions/surface concentration spatial correlations $\rho_{ec}$ for the GEM-MACH ensemble (left column), the EnsAI ensemble (middle column), and the static covariance (right column) at the location in North Carolina at 2015-06-05 18:00 (top row), 2015-06-07 18:00 (middle row), and 2015-06-09 18:00 (bottom row).
@@
As described in Section 8.1, to make the examination of the correlation matrix more tractable, we examine the different modes $m$ of the correlation length $L_{ec}[m]$ instead of dealing with the full correlation matrix directly. To illustrate the expansion into modes, Figure 13 shows the correlation length modes $L_{ec}[m]$ as a function of wavenumbers $m$ for the correlations in the top row of Fig. 12. In this example, the isotropic correlation lengths $L_{ec}[0]$ for the different covariances are between 65 and 90 km, while the anisotropic correlation length modes decrease rapidly with increasing wavenumber. Looking at the absolute value of the correlation length modes (top panel of Fig. 13), we can see that the non-isotropic modes for the static covariance are much smaller than that for the GEM-MACH and EnsAI ensembles, indicating less anisotropy in the static correlation. The EnsAI values for $L_{ec}[m]$ are within the 1–sigma confidence interval of the GEM-MACH value for many modes, and is generally much closer to the GEM-MACH values than the static covariance is. As $L_{ec}[m]$ decreases rapidly with increasing wavenumber, for the remainder of this paper we concentrate mainly on results for modes with $m\leq 2$ for brevity.
@@
Figure 13: Horizontal correlation length modes $L_{ec}[m]$ as a function of wavenumber $m$ for the emissions/surface concentration cross-correlation at the location in North Carolina at 2015-06-05 18:00. The top, middle, and bottom panels show the absolute, real, and imaginary parts, respectively.
@@
Time series of $L_{ec}[m]$ for $m\leq 2$ for the five locations identified in Fig. 1 are shown in the bottom four panels of Figs. 5 to 9. In all of these plots, $L_{ec}[m]$ for the EnsAI ensemble are much closer to the GEM-MACH ensemble value than the static covariance is and captures most of its time variation. While a diurnal pattern is evident in the GEM-MACH ensemble, there is significant non-diural time variation throughout the week that is dependent on meteorological conditions. In Fig. 5, we can see that the EnsAI correlation length modes generally follow the pattern in the original GEM-MACH ensemble, but the results are even more striking for the other four locations (Figs. 6 to 9), where the differences between the EnsAI and GEM-MACH ensembles are often smaller (or not much larger) than the sample variation in the GEM-MACH ensemble, or in other words, the differences between these values are statistically insignificant throughout much of the evaluation time period.
The three lower panels of Fig. 10 show the $ACC$ values for $L_{ec}[m\leq 2]$. For all months, EnsAI has much higher $ACC$ values than the static covariances. The $ACC(L_{ec}[m])$ values with $m\leq 2$ for the static covariances are close to zero for all cases, while the EnsAI values are above 0.5 for all cases with the exception of the $m=1$ mode in January and February, and are at or above 0.8 for $m=1,2$ for March to December. From this we can conclude that the EnsAI ensemble can accurately reproduce the time variations in the horizontal correlations (for modes with $m\leq 2$) that are not captured by the diurnal pattern of the static covariances. Similar results are seen for the $ACC$ values for $L_{cc}[m\leq 2]$ (see Figure S4 of the Supplement), although the $ACC(L_{cc}[1]/L_{cc}[0])$ values are closer to 0.6 for most months, in contrast to $ACC(L_{ec}[1]/L_{ec}[0])$ which is closer to 0.8.
EnsAI had lower $RMSE$ values than the static covariances for all months, as seen in the three lower panels of Fig. 11. For July to December, $RMSE(L_{ec}[0])$ for EnsAI is about 7 km less than the $RMSE$ value for the static covariances, which corresponds to about a 10% larger $RMSE$ value for the static covariances (see Figure S2 of the Supplement for the mean values of $L_{ec}[0]$). In some areas of the model domain during the summer and fall, $RMSE(L_{ec}[0])$ for the static covariances are over 25 km larger than the values for EnsAI (see Figure S5 of the Supplement), which correspond to static covariance $RMSE$ values which are between 50% to 490% larger than for EnsAI. The $RMSE$ values for $L_{ec}[1]/L_{ec}[0]$ taken over the whole domain are 0.07 to 0.14 less for EnsAI than for the static covariance, and between 0.04 and 0.09 less for EnsAI for $L_{ec}[2]/L_{ec}[0]$, as displayed in the two lowest panels of Fig. 11. Additionally, there are many locations within the model domain where $RMSE(L_{ec}[1]/L_{ec}[0])$ and $RMSE(L_{ec}[2]/L_{ec}[0])$ for the static covariances are over 0.3 and 0.2 higher, respectively, as compared to the $RMSE$ value for EnsAI (see Figures S6 and S7 of the Supplement). Lastly, we note that the EnsAI $RMSE$ values for the surface concentration univariate correlation length modes $L_{cc}[m]$ are uniformly less than those for the static covariances for $m\leq 2$ (see Figure S8 of the Supplement).
@@
To get an intuitive understanding of what these differences in correlation length modes correspond to in terms of spatial patterns, Figure 14 shows another set of example correlations that illustrate the deviations in correlation length mode values mentioned in the previous paragraph. The leftmost panel of Fig. 14 shows an isotropic correlation with $L[0]=65$ km. For illustrative purposes, we will think of this isotropic correlation as a target correlation (i.e. from the GEM-MACH ensemble) and compare it to correlations with different $L[m]$ values to see the difference in the spatial patterns. The correlation shown in the panel second from the left has $L[m\leq 2]$ values that deviate from the isotropic correlation in the leftmost panel by amounts representative of the $RMSE$ values found in Fig. 11 for EnsAI. This correlation is only slightly anisotropic and slightly smaller than the correlation in the leftmost panel. The correlation in the panel second from the right has values that deviate from the isotropic correlation close to the $RMSE$ values for the static covariance in Fig. 11 and is noticeably more anisotropic than the two correlations to the left. The correlation in the rightmost panel has values that deviate from the isotropic correlation by amounts equal to differences between the GEM-MACH and static covariance values that are commonly found in the model domain (see Figures S6 and S7 of the Supplement), which produces a correlation with a drastically different spatial pattern than the correlation in the leftmost panel.